40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 2 11:37:38 2019, gautam, Update, ALS, EY IR ALS Assay IMG_7780.JPGEy_MM_20190802.pdf
    Reply  Tue Aug 6 23:09:20 2019, gautam, Update, ALS, EY table work 
       Reply  Fri Aug 9 16:37:39 2019, gautam, Update, ALS, More EY table work IMG_7791.JPGfiberCleaning.pdf
Message ID: 14835     Entry time: Tue Aug 6 23:09:20 2019     In reply to: 14823     Reply to this: 14838
Author: gautam 
Type: Update 
Category: ALS 
Subject: EY table work 
  • Removed power monitoring PD (It was off anyways)
  • Installed Steering mirror and collimator in K6XS mount (fast axis = p-pol to best effort by eye)
  • Installed lens mounts in approx position
  • Cleaned fiber at EY and connected to the collimator
  • Coupled EY--->PSL and spare PSL-->EY fibers together at the PSL table to facilitate coupling.
  • tbc tomorrow...
Quote:

Couple IR light into fiber with good MM at EY

ELOG V3.1.3-